OpenStudy (anonymous):

What is the slope of a line parallel to the line whose equation is 5x + 3y = 15?

OpenStudy (math&ing001):

First we find the slope of 5x + 3y = 15

OpenStudy (anonymous):

i dont know how to

OpenStudy (anonymous):

ive always been confused with slopes

OpenStudy (math&ing001):

Do it like you're trying to solve the equation for y.

OpenStudy (anonymous):

how bout you solve for y first might i suggest

OpenStudy (anonymous):

then what ever is times x is your slope

OpenStudy (math&ing001):

Example 4y+x = 6 ie y = -1/4 x + 6/4 Here -1/4 is the slope

OpenStudy (anonymous):

that makes a little bit more sense

OpenStudy (math&ing001):

Well can you do it now ?

OpenStudy (anonymous):

ill try to find the slope

OpenStudy (anonymous):

is the slope 5/3?

OpenStudy (anonymous):

remember solve for y and what ever is multiplying x is your slope

OpenStudy (math&ing001):

You forgot the (-) sign !

OpenStudy (anonymous):

-5/3?

OpenStudy (anonymous):

it would be negative but yes and your y intercept would be 5

OpenStudy (anonymous):

correct

OpenStudy (math&ing001):

Yep ! So now each line that has a slope of -5/3 is parallel to the line whose equation is 5x + 3y = 15

OpenStudy (anonymous):

ok...thank you!!!

OpenStudy (anonymous):

welcome

OpenStudy (math&ing001):

YW